Bill and Janine both changed jobs recently and each received a rollover check for $5,000. Both would like to invest in the ABC Mutual Fund (historical prices in the table below.)

• Bill decided to invest all of his money on June 1st.
• Janine decided to dollar cost average her funds into the mutual fund by investing $1,000 on the first of the month for five months starting June 1st, continuing through October 1st .

A. Use the following table to determine the value of both of their accounts as of November 1st .

B. Show how many shares were purchased for each investment (one for Bill and five for Janine).

C. Explain why the two account balances differ.

ABC Mutual Fund

Date
June 1
July 1 August 1 September 1 October 1 November 1

Price 10.21 9.97 10.22

9.67 10.13 10.50
Answer #1

Answer:

Given that

Each received a rollover check for $5,000

ABC mutual fund

June 1     10.21

August1    9.97

September1 10.22

October1       9.67

November1   10.13

Price            10.50

Use the following table to determine the value of both of their accounts as of November 1st

Answer (A):

Bill

Janine

Date

Price (p)

Amount invested (a)

Shares bought (s = a/p)

Value (s*p)

Amount invested (a)

Shares bought (s = a/p)

Value (s*p)

June 1

10.21

5,000.00

(5,000.00/10.21) = 489.72

1,000.00

97.94

July 1

9.97

1,000.00

100.30

August 1

10.22

1,000.00

97.85

September 1

9.67

1,000.00

103.41

October 1

10.13

1,000.00

98.72

November 1

10.50

(489.72*10.50)= 5,142.02

5,231.32

Total

489.72

498.22

A). Value of Bill's account on November 1st is $5,142.02

Value of Janine's account on November 1st is $5,231.32

Answer (B): As seen from the table above, Bill bought 489.72 or 490 offers and Janine acquired 498.22 or 498 offers.

Answer(C). The two record adjusts vary on the grounds that Bill puts the whole $5,000 in one go while Janine puts resources into month to month openings of $1,000. This implies she pays an alternate cost for each $1,000 opening she purchases. This is the motivation behind why both records adjusts contrast,